OpenStudy (anonymous):

A pharmacist wishes to mix a solution that is 6% Minoxidil. She has on hand 100ml of a 1% solution and wishes to add some 8% solution to obtain the desired 6% solution. How much 8% solution should she add?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I think the answer is 250ml but I'm really not sure and I need help please :)

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (kropot72):

The amount of Minoxidil in 100ml of 1% solution = 100 * 0.01 = 1ml. Let the required amount of 8% solution be x ml. Now we can write the following equation based on amounts of Minoxidil: \[\large (100\times0.01)+0.08x=0.06(100+x)\ ..........(1)\] Simplifying (1) gives: \[\large 1+0.08x=6+0.06x\ .........(2)\] Now you need to solve (2) to find x, the number of mls of 8% solution that is needed.